OpenStudy (anonymous):

whats the replacement for x 2x-36=-44

OpenStudy (anonymous):

2x = -44 + 36 2x = -8 Can you find x?

OpenStudy (anonymous):

2x=36-44 x=(36-44)/2

OpenStudy (anonymous):

is the answer -4??

mathslover (mathslover):

good michael